A triode passes a plate current of 5mA at plate voltage 150V and grid voltage of -2V. If the grid voltage is changed to -35V, the plate current drops to 3.2mA but can be restored to 5mA by increasing plate voltage to 195V. Calculate


(1) mutual conductance


(2) a.c. plate resistance


(3) amplification factor


1
Expert's answer
2022-02-14T14:28:00-0500

1) mutual conductance

"g_m= \frac{change\> in \>anode\> current}{change\>in\> grid\> voltage}"


"=\frac{(3.2-5)×10^{-3}}{-35-(-2)}=5.4645×10^{-5}mho"


2) ac plate resistance

"r_p=\frac{small\> increment\> in \>anode\> voltage}{Small\> change\, in\> anode\> current}"


"=\frac{195-150}{(5-3.2)×10^{-3}}=2.5×10^4\Omega"



3) Amplification factor


"\mu=\frac{change\> in\> anode \>voltage}{change\> in \>grid \>voltage}"


"=\frac{195-150}{-35-(-2)}=-1.3636"